OpenStudy (anonymous):

A bag contains 4 green, 4 red, and 2 blue crayons. What is the probability of randomly choosing 2 red crayons without replacement. write your answer as a decimal

OpenStudy (anonymous):

4/10 * 3/9 = 2/15= 0.13 Hoped this helped :)